Exactly 500 cubic centimeters of a gas at 1 atmosphere and 0 degrees Celsius weighs 0.581 grams. The composition of the gas shows it to be composed of only carbon and hydrogen, the carbon being 92.24%. What is the molecular formula of this gas?
1
Expert's answer
2011-10-27T12:56:37-0400
The STP are mentioned -> we can use the law, that 1 mole of any gas has 22,4L
nu=V/V0
nu=0.5/22.4=0.02232 mol
M=m/nu
M=0.581/0.02232=26
omega=M(C)*x/(M(C)*x+M(H)*y)
0.9224=12x/(12x+y)
11.0688x+0.9224y=12x
0.9312x=0.9224y
x is approximately equal to y
When x=1 y=1 M=12+1=13 no
& x=2 y=2 M=12+12+1+1=26 YES!
& x=3 y=3 M= 12+12+12 +1+1+1=39 too much

Answer: C2H2